# Review of multi-fidelity models

### v5

@@ 5.3. Cost savings and accuracy report
Property Value
a Comments
Cost LF/HF
0.07
a LF= Euler, HF= RANS
Error LF/HF
0.18
Cost MF/HF
0.13
a MF= 1 HF + 17 LF
Error MF/HF
0.05
@@ Appendix E Tables
Dimensionality
[68] 2D/3D Eu, [101] 1D/3D RANS+TM, [105] 2D/3D URANS, [126] 2D/3D, [174] 1D/2D RANS, [189] 1D/2D Li, [214] 1D/3D RANS, [228] 1D/3D RANS, [247] 1D,2D/3D RANS
Coarse/Refined
[5] Eu, [30] RANS, [41] Eu, [42] Eu, [64] Eu, [43] Li/Eu, [102] RANS, [108] MPF, [111] MHD, [117] Eu, [118], Eu[121] Eu, [134] Eu, [141] RANS, [184] RANS, [209] RANS, [237] Eu/RANS
Exp./Sim.
[65] Euler/MHD, [67] PF/Em, [123] RANS, [210] RANS
Semiconverged/Converged
[102], RANS[118] Eu
Steady/Transient
[24] AE, [75] Eu, [209] TM
@@
Dimensionality
[130] 1D/2D Li, [129] 1D/3D, [145] 2D/3D, [146] 2D/3D Li, [199] 2D/3D Li, [201] 2D/3D NL
Coarse/Refined
[17] Li, [25] NL, [27] Li, [29] NL, [37] Li, [132] Li, [144] Li, [202] NL, [203] NL, [226] Li, [235] Li, [236] Li
Boundary conditions
[222] Li, [225] Li
@@
Additive correction
[17] [16] [32] [58] [81] [113] [164] [188] [190] [193] [198] [199] [203] [202] [213] [221] [225]
Multiplicative correction
[4] [5][17] [16] [31] [32] [37] [81] [86] [92] [97] [106] [141] [146] [157] [188] [190] [198] [199] [203] [202] [213] [218] [219] [222] [224] [225] [226]
Comprehensive correction
[59] [76] [110] [159] [191] [235] [236]
Space mapping
[34] [102] [121] [183] [189]
@@
Additive correction
[25] [68] [89] [111] [143] [144] [169] [174] [179]
Multiplicative correction
[40] [67] [142]
Comprehensive correction
[8] [29] [30][71] [80] [96] [107] [108] [123] [128] [127] [130] [129] [133] [172] [208] [210] [234]
Calibration + comprehensive correction
[23] [65] [93] [109]

### v6

@@ 5.3. Cost savings and accuracy report
Property Value a Comments
Cost LF/HF 0.07 a LF= Euler, HF= RANS
Error LF/HF 0.18 -
Cost MF/HF 0.13 a MF= 1 HF + 17 LF
Error MF/HF 0.05 -
@@ Appendix E Tables
Dimensionality [68] 2D/3D Eu, [101] 1D/3D RANS+TM, [105] 2D/3D URANS, [126] 2D/3D, [174] 1D/2D RANS, [189] 1D/2D Li, [214] 1D/3D RANS, [228] 1D/3D RANS, [247] 1D,2D/3D RANS
Coarse/Refined [5] Eu, [30] RANS, [41] Eu, [42] Eu, [64] Eu, [43] Li/Eu, [102] RANS, [108] MPF, [111] MHD, [117] Eu, [118], Eu[121] Eu, [134] Eu, [141] RANS, [184] RANS, [209] RANS, [237] Eu/RANS
Exp./Sim. [65] Euler/MHD, [67] PF/Em, [123] RANS, [210] RANS
Semiconverged/Converged [102], RANS[118] Eu
Steady/Transient [24] AE, [75] Eu, [209] TM
@@
Dimensionality [130] 1D/2D Li, [129] 1D/3D, [145] 2D/3D, [146] 2D/3D Li, [199] 2D/3D Li, [201] 2D/3D NL
Coarse/Refined [17] Li, [25] NL, [27] Li, [29] NL, [37] Li, [132] Li, [144] Li, [202] NL, [203] NL, [226] Li, [235] Li, [236] Li
Boundary conditions [222] Li, [225] Li
@@
Additive correction [17] [16] [32] [58] [81] [113] [164] [188] [190] [193] [198] [199] [203] [202] [213] [221] [225]
Multiplicative correction [4] [5][17] [16] [31] [32] [37] [81] [86] [92] [97] [106] [141] [146] [157] [188] [190] [198] [199] [203] [202] [213] [218] [219] [222] [224] [225] [226]
Comprehensive correction [59] [76] [110] [159] [191] [235] [236]
Space mapping [34] [102] [121] [183] [189]
@@
Additive correction [25] [68] [89] [111] [143] [144] [169] [174] [179]
Multiplicative correction [40] [67] [142]
Comprehensive correction [8] [29] [30][71] [80] [96] [107] [108] [123] [128] [127] [130] [129] [133] [172] [208] [210] [234]
Calibration + comprehensive correction [23] [65] [93] [109]